# Interpretability of linear regression models of glassy dynamics

### v1

@@ VI.2.3 Linear regression in the PCA basis
Let us first focus on the OLS regression case. As shown in the Appendix D, the dynamic propensity predicted by OLS regression is
@@ VI.2.4 Dimensional reduction and comparison with elastic net
Figure 9: Pearson coefficient $R$ between the ground truth dynamic propensity and PCR predictions using the first $P$ PCs sorted according to their eigenvalue (squares) or their Pearson coefficient with the dynamic propensity (circles). The results of the optimal elastic net regression models are also included (triangles). Note the change of scale on the x-axis after $P=5$ components. The tics on the right side of the figure indicate the values of the Pearson coefficient reached for a given value of number of PCs included in the regression model.
It is interesting to compare these results to those obtained by the optimal elastic net models, which are included as dashed lines in Fig. 9. We see that the elastic net models converge to the maximum correlation, as a function of selected features, approximately like in supervised PCR. For small numbers of features, the two regression models provide about the same performance accuracy, although the precise details depend on the descriptor. One advantage of PCR is that the feature extraction scheme is straightforward and efficient, thanks to the orthogonality of the PCA basis. By contrast, although the features selected by elastic net are not orthogonal to each other, they are more easily interpretable than the PC eigenvectors. The latter, in fact, mix all the features of the original descriptor and do not always lean themselves to a straightforward interpretation. We will further discuss this issue in Sec. VII.
@@
Figure 10: Dynamic propensity of a representative configuration: (a) ground truth and estimates from (b) the OLS regression model, (c) the PCR model with $P=5$ components, and (d) the PCR model with $P=2$ components. Dark and light particles correspond to fast and slow particles, respectively. The mid panels, from (e) to (h) show the same as in the top panels coarse-grained over a length $\ell=1.5$. Particles are colored in black and white according to whether the corresponding coarse-grained variable is above or below the average, respectively. (i) Probability density function of the ground truth and estimated dynamic propensities for various models, using all the available configurations. The remaining panels in the bottom row display scatter plots of the ground truth propensity against its estimated value for (j) the OLS regression model, (k) the PCR model with $P=5$ components, and (l) the PCR model with $P=2$ components. In all the panels, $\hat{Y}$ is shifted and scaled to match the mean and variance of the ground truth $Y$.
@@ VII Discussion
Real space structure of the models’ prediction– A first piece of information comes from direct visual comparison of the ground truth dynamic propensities and the predictions of the linear regression models. Here, we consider two low-dimensional PCR models, namely the ones with $P=2$ and $P=5$ within supervised PCR, as well as with the full $P=M$ model (corresponding to the OLS regression model) obtained with the BP descriptor. In this section, we revert our initial feature normalization (see Sec. III.5): we scale and shift $\hat{Y}_{i}$ obtained from the PCR models so that their mean and variance match those of the actual unscaled dynamic propensity. Thus, we will consider $\hat{Y}_{i}$ vs $Y_{i}$ in their original LJ units.
Figure 10 shows the distributions of $\hat{Y}_{i}$ vs $Y_{i}$, the scatter plots of $\hat{Y}_{i}$ vs $Y_{i}$ and the corresponding snapshots for a representative configuration in our dataset. We see that the full descriptor ($P=M$ or OLS regression model) provides an excellent description of the dynamic propensity field, which is accurately reconstructed in most of its details, except for a slight discrepancy in the shape of the distribution, see Fig. 10(i). Thus, $R$ values of the order of 0.9 correspond to a very satisfactory description of the dynamic propensity. The model obtained with the $P=5$ most correlated PCs gives $R\approx 0.7$ and reproduces most of the patterns of the dynamic propensity field, despite some additional noise. The correspondence is good when considering coarse-grained fields (see the mid panels), indicating that such models are able to grasp the relevant dynamic fluctuations on large length scales [75]. The model with the $P=2$ most correlated PCs, instead, only captures some of the fast and slow regions of the propensity, but the overall large-scale structure of the propensity field is not accurately reproduced. Qualitatively, these results suggest that correlation coefficients of 0.55, 0.7, 0.9 correspond to modest, satisfactory, and excellent estimation of the dynamic propensity, respectively.
@@
Even when using inherent structures, however, a clear difference can be seen at the level of parsimonious models. Interestingly, the low-dimensional models obtained from PCR perform better with the SLO descriptor than with the BP and JBB descriptors. This might be explained by the inclusion from the outset of some physically relevant features like the $\Theta$ parameter. With the SLO descriptor, in fact, even a two-features model achieves a high Pearson coefficient ($R\approx 0.81$) and very little is gained by including the full set of features ($R\approx 0.86$). By contrast, high-dimensional descriptors accurately reconstruct the observable of interest, exploiting by brute force a broad spectrum of features, but they may fail to provide an interpretable parsimonious models. This could also partly depend on the details of the descriptor: the atomic cluster expansion, for instance, successfully predicts plasticity in amorphous solids even upon strong dimensional reduction [76].
@@
To sum up on this point, PCR of the SLO descriptor yields a simple two-state structural model. The two relevant variables account for steric effects on an intermediate length scale and for some of the fluctuations of bond-orientational order. Superficially, these results suggest a connection with the phenomenological two-state model developed by Tanaka in the 1990s [53], which included local density and bond-order fluctuations as structural order parameters. We emphasize, however, that our ternary model does not display strong 6-fold orientational order and $\overline{\Psi}_{6}$ itself is poorly correlated to the dynamic propensity. It is only when the fluctuations of $\overline{\Psi}_{6}$ are coupled to those of the other features as described by $\textbf{u}^{(5)}$ that some correlation emerges. We found that the other descriptors consistently identify a structural mode related to local packing (e.g., PC2 in the BP descriptor, PC1 in the JBB descriptor) that is moderately correlated to the dynamics, but none that is specifically connected to bond-orientational order.
@@
Optimality criteria– Are there quantiative criteria to decide whether to accept a low-dimensional model, which is easier to interpret, at the expense of performance? We considered several optimality criteria, developed in the context of statistical learning, that attempt to define an optimum model by balancing the performance and the model complexity. In particular, we employed two well-known scores: the Akaike Information Criterion (AIC) [77] and the Bayesian Information Criterion (BIC) [78]. Both criteria address over-fitting by combining the maximized log-likelihood, which quantifies data fidelity, with a complexity penalty term that grows with the number of parameters. In general, AIC applies a lighter penalty and therefore tends to keep more features if they help prediction, while BIC is stricter and usually picks the smaller model that still describes the data well. We evaluated AIC and BIC for our glass-forming liquids dataset using Lasso regression with varying magnitudes of the L1 penalty, which controls the number of selected features. Both AIC and BIC indicated that the optimal model corresponds to OLS (i.e., including all features). Thus, in our setting, AIC and BIC do not provide any practical guidance.
@@
Our current standpoint on this issue is that there is at present no robust quantitative criterion to select an optimal model that strikes a balance between prediction and model size. We think that the simple empirical criterion of selecting a small number of features, say up to 5, is in line with a well-established approach when building phenomenological models in statistical physics [51]. We suggest that looking at the best two- or five-feature model provides a basic “handful criterion” to identify parsimonious data-driven models of glassy dynamics.
@@ VIII Conclusions
To identify low-dimensional linear models of glassy dynamics, we used two simple dimensional reduction techniques. First, we considered elastic net regression, which yields a set of parsimonious models that strike a good balance between accuracy and physical interpretability. Second, we performed principal component regression using a supervised selection scheme of the principal components. This approach yields an accurate enough description of the dynamic propensity with a handful of collective structural features. Overall, our work establishes that linear regression models, once properly fine-tuned, can be useful tools to identify the relevant structural modes associated to dynamic heterogeneities in glass-forming liquids. It would be interesting to extend our analysis to more sophisticated high-dimensional descriptors and to other glassy systems beyond the simple two-dimensional model liquid considered here, including amorphous solids subjected to external loading.
@@ Acknowledgements.
The data and workflow necessary to reproduce the findings of this study will be openly available after publication of the paper in the Zenodo data repository.
Appendix A Selected results for the SLO and JBB descriptors
@@
In this section, we present additional results for the physically motivated descriptors introduced in Sec. III.4. Figures 14 and 14 display the Pearson coefficients $R[{X^{(f)}},{Y}]$ between the dynamic propensity and each structural feature $\textbf{X}^{(f)}$ of the SLO and JBB descriptors, respectively.
@@
It is instead more difficult to extract straightforward physical information from the analysis of the JBB descriptor, due to its complex dependence on the chemical composition. It is nonetheless interesting to note that the perimeter $\overline{\pi}$ of the Voronoi cells, which is included in the JBB descriptor but not in the SLO one, has an overall negative correlation with the dynamic propensity. This again points to a coupling between steric effects, i.e., larger local free volume, and mobility.
@@
In Figures 16 and 16, we show the correlation matrices $\mathrm{C}$ of the SLO and JBB descriptors, respectively. As for the BP descriptor, the correlation matrices display a pronounced block structure: there are strong correlations within subblocks of features, this time due to similar coarse-graining lengths, as well as non-trivial correlations between different features. In particular, the SLO descriptor reveals a negative correlation between local number density $\overline{\rho}$ and local packing fraction $\overline{\varphi}$. This counter-intuitive result is likley a result of compositional fluctuations in this mixture, see Sec. III.6.
@@
Appendix B Ordinary least squares regression and Ridge regression

### v2

@@ VI.2.3 Linear regression in the PCA basis
Let us first focus on the OLS regression case. As shown in the Appendix E, the dynamic propensity predicted by OLS regression is
@@ VI.2.4 Dimensional reduction and comparison with elastic net
Figure 10: Pearson coefficient $R$ between the ground truth dynamic propensity and the PCR predictions using the first $P$ PCs sorted according to their eigenvalue (squares) or their Pearson coefficient with the dynamic propensity (circles). The results of the optimal elastic net regression models are also included (triangles). Note the change of scale on the $x$ axis after $P=5$ components. The tics on the right side of the figure indicate the values of the Pearson coefficient reached for a given value of number of PCs included in the regression model.
It is interesting to compare these results to those obtained by the optimal elastic net models, which are included as dashed lines in Fig. 10. We see that the elastic net models converge to the maximum correlation, as a function of selected features, approximately like in supervised PCR. For small numbers of features, the two regression models provide about the same performance accuracy, although the precise details depend on the descriptor. One advantage of PCR is that the feature extraction scheme is straightforward and efficient, thanks to the orthogonality of the PCA basis. By contrast, although the features selected by elastic net are not orthogonal to each other, they are more easily interpretable than the PC eigenvectors. The latter, in fact, mix all the features of the original descriptor and do not always lend themselves to a straightforward interpretation. We will further discuss this issue in Sec. VII.
@@
Figure 11: Dynamic propensity of a representative configuration. (a) Ground truth and estimates from (b) the OLS regression model, (c) the PCR model with $P=5$ components, and (d) the PCR model with $P=2$ components. Dark and light particles correspond to fast and slow particles, respectively. The mid panels, from (e) to (h) show the same as in the top panels coarse-grained over a length $\ell=1.5$. Particles are colored in black and white according to whether the corresponding coarse-grained variable is above or below the average, respectively. (i) Probability density function of the ground truth and estimated dynamic propensities for various models, using all the available configurations. The remaining panels in the bottom row display scatter plots of the dynamic propensity against its estimated value for (j) the OLS regression model, (k) the PCR model with $P=5$ components, and (l) the PCR model with $P=2$ components. In all the panels, $\hat{Y}$ is shifted and scaled to match the mean and variance of the dynamic propensity pp.
@@ VII Discussion
Real space structure of the models’ prediction– A first piece of information comes from direct visual comparison of the ground truth dynamic propensities and the predictions of the linear regression models. Here, we consider two low-dimensional PCR models, namely the ones with $P=2$ and $P=5$ within supervised PCR, as well as with the full $P=M$ model (corresponding to the OLS regression model) obtained with the BP descriptor. In this section, we revert our initial feature normalization (see Sec. III.5): we scale and shift the predicted propensities, $\hat{Y}_{i}$ so that their mean and variance match those of the actual unscaled dynamic propensity. Thus, we will consider $\hat{Y}_{i}$ vs. $p_{i}$ in the original LJ units.
Figure 11 shows the distributions of $\hat{Y}_{i}$ vs $p_{i}$, the scatter plots of $\hat{Y}_{i}$ vs $p_{i}$ and the corresponding snapshots for a representative configuration in our dataset. We see that the full descriptor ($P=M$ or OLS regression model) provides an excellent description of the dynamic propensity field, which is accurately reconstructed in most of its details, except for a slight discrepancy in the shape of the distribution, see Fig. 11(i). Thus, $R$ values of the order of 0.9 correspond to a very satisfactory description of the dynamic propensity. The model obtained with the $P=5$ most correlated PCs gives $R\approx 0.7$ and reproduces most of the patterns of the dynamic propensity field, despite some additional noise. The correspondence is good when considering coarse-grained fields (see the mid panels), indicating that such models are able to grasp the relevant dynamic fluctuations on large length scales [75]. The model with the $P=2$ most correlated PCs, instead, only captures some of the fast and slow regions of the propensity, but the overall large-scale structure of the propensity field is not accurately reproduced. Qualitatively, these results suggest that correlation coefficients of 0.55, 0.7, 0.9 correspond to modest, satisfactory, and excellent estimation of the dynamic propensity, respectively.
@@
Even when using inherent structures, however, a clear difference can be seen at the level of low-dimensional models: the low-dimensional models obtained from PCR using the SLO descriptor perform better than those obtained using the BP and JBB descriptors. This might be explained by the inclusion from the outset of some physically relevant features like the $\Theta$ parameter. With the SLO descriptor, in fact, even a two-features model achieves a high Pearson coefficient ($R\approx 0.81$) and very little is gained by including the full set of features ($R\approx 0.86$). By contrast, high-dimensional descriptors accurately reconstruct the observable of interest, exploiting by brute force a broad spectrum of features, but they may fail to provide a simple parsimonious model. This could also partly depend on the details of the descriptor: the atomic cluster expansion, for instance, successfully predicts plasticity in amorphous solids even upon strong dimensional reduction [76].
@@
To sum up on this point, PCR of the SLO descriptor yields a simple two-feature structural model. The two relevant variables account for steric effects on an intermediate length scale and for some of the fluctuations of bond-orientational order. Superficially, these results suggest a connection with the phenomenological two-state model developed by Tanaka in the 1990s [77], which included local density and bond-order fluctuations as structural order parameters. We emphasize, however, that our ternary model does not display strong 6-fold orientational order and $\overline{\Psi}_{6}$ itself is poorly correlated to the dynamic propensity. It is only when the fluctuations of $\overline{\Psi}_{6}$ are coupled to those of the other features as described by $\textbf{u}^{(5)}$ that some correlation emerges. We found that the other descriptors consistently identify a structural mode related to local packing (e.g., PC2 in the BP descriptor, PC1 in the JBB descriptor) that is moderately correlated to the dynamics, but none that is specifically connected to bond-orientational order.
Figure 14: Sequence of sorted weights $w$ as a function of sorted feature index $f$ in (a) Ridge regression with $\alpha=0.1$ and (b) PCR. Note that in PCR the weights coincide with the Pearson coefficients $R[{\tilde{X}},{Y}]$, see Eq. (48).
Feature selection– One practical approach to feature selection is to look for gaps in the weight distribution of the linear models: features whose absolute weights stand out from the bulk are then identified as the most relevant ones. In Fig. 14, we illustrate these ideas for Ridge regression with $\alpha=0.1$ and PCR using the BP descriptor. In each panel, we show the sequence of sorted weights $w$ as a function of sorted feature index $f$. By visual inspection, we see that no more than four or five features show weights that stand out from the rest in such an ordered sequence. A preliminary analysis of the weight distributions support this idea. In Ridge regression, the relevant features pinpoint structural modes involving radial features localized around the first peaks of the partial radial distribution functions. In PCR, the principal components from PC2 to PC5 stand out, as was also clear from Fig. 8. Interestingly, we found that the two features with largest absolute weights in Ridge regression, i.e., $G^{S}(8)$ and $G^{L}(10)$, have similar projections on the physically motivated variables as PC1 and PC2, respectively. These weights’ sequences provide good examples of sparse representations that support physical interpretation.
@@
Optimality criteria– Are there quantitative criteria to decide whether to accept a low-dimensional model, which is easier to interpret, at the expense of performance? We considered several optimality criteria, developed in the context of statistical learning, that attempt to define an optimum model by balancing the performance and the model complexity. In particular, we employed two well-known scores: the Akaike information criterion [78] and the Bayesian information criterion [79]. Both criteria address over-fitting by combining the maximized log-likelihood, which quantifies data fidelity, with a complexity penalty term that grows with the number of parameters. We found, however, that these criteria do not provide any practical guidance in our setting, since we operate in a high-data regime where overfitting is negligible.
@@
Our current standpoint on this issue is that at present there is no robust quantitative criterion to select an optimal model that strikes a balance between prediction and model size. Selecting a small number of physical variables, say up to five, is in line with a well-established approach to develop phenomenological models in statistical physics [53]. We thus suggest that a parsimonious data-driven description of glassy dynamics should involve no more than a handful of relevant features.
Dependence on spatial dimensions– Characterizing the local structure of glassy systems may appear an easier task in 2D than in 3D. Provided the input structural descriptor is complete enough, however, linear regression and deep learning models of glassy dynamics eventually achieve similar performances in both spatial dimensions [24]. While 3D systems may pose a more serious challenge for dimensional reduction, the same kind of physically motivated features we used in this study as explanatory variables can be readily applied to 3D systems. Therefore, at least from the methodological viewpoint, there is no qualitative difference between studying 2D and 3D systems.
Figure 15: Cross-state generalization of the propensity of motion $p(t=\tau_{\alpha};T)$ from Ridge regression at a reference temperature $T_{r}$. The solid line indicates the in-state prediction (interpolation), while the thin lines with symbols indicate the cross-state predictions (extrapolation), from $T_{r}=0.30$ (circles) and $T_{r}=0.40$ (squares).
Cross-state generalization– Throughout this work we focused on the performance of linear regression at the same temperature at which the model was trained. As anticipated in Sec. II, from a physicist’s viewpoint, data interpolation does not qualify as a genuine form of prediction. To assess the actual predictive power of Ridge regression, we show in Fig. 15 the results for the cross-state generalization of the propensity of motion at $t=\tau_{\alpha}$: the model’s weights are fitted at a reference temperature $T_{r}$ using $\alpha=0.1$ and then used to extrapolate the propensity of motion at a different temperature $T$. The extrapolation from $T_{r}=0.30$ to higher temperatures provides good predictions up to around the onset temperature $T\approx 0.5$, where the Pearson coefficients start to decrease significantly even for in-state predictions. Interestingly, the model trained at $T_{r}=0.40$ provides consistent predictions over a broad range of temperatures, covering around three decades in structural relaxation times [54]. These findings resemble those of obtained in Ref. 24 using linear regression and deep learning models. Note, however, that all these ML models can only predict the propensity of motion at a time $t=\tau_{\alpha}$, i.e., the absolute time scale must be determined independently. Thus, current ML models are not fully predictive yet.
@@ VIII Conclusions
To identify low-dimensional linear models of glassy dynamics, we used two simple dimensional reduction techniques. First, we considered elastic net regression, which yields a set of low-dimensional models that strike a good balance between accuracy and physical interpretability. Second, we performed principal component regression using a supervised selection scheme of the principal components. This approach yields an accurate enough description of the dynamic propensity with a handful of collective structural features. Overall, our work establishes that linear regression models, once properly fine-tuned, can be useful tools to identify the relevant structural modes associated to dynamic heterogeneities in glass-forming liquids. It would be interesting to extend our analysis to more sophisticated high-dimensional descriptors and to other glassy systems beyond the simple two-dimensional model liquid considered here, including molecular and polymeric liquids as well as amorphous solids subject to external loading.
@@ Acknowledgements.
The data and workflow necessary to reproduce the findings of this study are available in the Zenodo data repository [80].
Appendix A Species-resolved analysis
In multi-component mixtures, the particle mobility is in general species dependent. Therefore, most studies of structure-dynamics relationship analyze separately each species, or even just focus on the majority species. In the body of the text, we chose to analyze all the particles irrespective of their species. As we will show in this appendix, the key findings of our work do not depend on whether the analysis is performed species-wise or for the full system.
In Fig. 16 we show the probability density functions for the dynamic propensity for particles of species S, M, and L. The distribution functions depend of course on the species, although the dependence is not dramatic. We found that the average values and corresponding standard deviations of the distributions are $0.75\pm 0.47$, $0.58\pm 0.37$, and $0.43\pm 0.25$ for species S, M, and L, respectively. For comparison, the corresponding result for the dynamic propensity of all the particles is $0.60\pm 0.41$. Since we could not find detailed results regarding species-dependent propensities in the literature [41], we cannot determine whether the above differences are smaller than those for other 3d glass-forming liquids.
Figure 16: Probability density function of the dynamic propensity for particles of species S (blue), M (red), and L (orange).
Figure 17: Species-wise (full lines) and total (dashed lines) prediction performance metrics for Ridge regression of the dynamic propensity using the BP descriptor: (a) Pearson coefficient $R[Y,\hat{Y}$] as a function of $\alpha$ and (b) coefficient of determination $R_{2}[Y,\hat{Y}$] as a function of $\alpha$.
One might expect that analyzing structure-dynamics correlations for the entire system will yield a different performance compared to a species-wise analysis. As shown in Fig. 17, this is indeed the case, but the effect is very small. In particular, for the optimal regularization parameter $\alpha=0.1$, we obtained Pearson coefficients between the BP descriptor and the propensity of 0.86, 0.85, 0.84 for species S, M, and L, respectively. These values can be compared with $R=0.85$ obtained for the full system, see Table. 3. Somewhat larger discrepancies are observed at very small $\alpha$, or in the absence of regularization, for the minority species (M). Overall, however, the performance of Ridge regression displays the same qualitative trends when performed species-wise or for the full system. We also confirm the presence of oscillations in the weights for $\alpha=0$ in the species-wise analysis.
Figure 18: Pearson coefficients $R$ between the first five PCs and physically motivated features coarse-grained over a length $\ell=1.5$. The analysis is performed species-wise for (a) species S, (b) species M, and (c) species L.
We now turn our attention to the PCA of the BP descriptor. Overall, the trend of PCA is qualitatively similar when performed species-wise or for the full system. The observation that PC1 is poorly correlated to the dynamic propensity is confirmed for all species. We found that PC2 is again the most correlated for species S and M, as for the full system. For species L, it is PC3 to be most correlated. These results suggest that it may be difficult to extract robust information from PCA beyond the very first few PCs (in $\lambda$ order).
Interestingly, we found that also the physical interpretation of the first two PCs is essentially the same across species. In Fig. 18, we show at the Pearson coefficients between the first 5 PCs of the BP descriptor and the physically motivated features. The latter features are coarse-grained over $\ell=1.5$ and the analysis is performed species-wise in each panel. The key structure of the correlation matrix for the species-resolved analysis is similar to the one of full system. In fact, PC1 and PC2 are projected similarly on the physically resolved features. Note, however, that for species L it is PC3 having the largest correlation with the dynamics and not PC2. It is difficult to grasp a clear interpretation of components of order higher than two in terms of physically motivated features.
Figure 19: Species-wise (full lines) and total (dashed lines) Pearson coefficient $R$ between the ground truth dynamic propensity and the PCR predictions using the first $P$ PCs sorted according to their eigenvalue (squares). Note the change of scale on the x-axis after $P=5$ components.
Finally, we assess the species dependence of the performance of PCR. Figure 19 shows the Pearson coefficient $R$ in PCR as function of the number $P$ of included collective features. The latter are sorted according to their individual Pearson coefficients. The growth of the Pearson coefficient as a function of $P$ is qualitatively similar for all the species and closely tracks the one for the total system. We conclude that although the details about the PCs with small eigenvalues and small correlation will depend in general of the selected species, the general trends discussed in the main text are robust.
Appendix B Selected results for the SLO and JBB descriptors
@@
In this appendix, we present additional results for the physically motivated descriptors introduced in Sec. III.4. Figures 21 and 21 display the Pearson coefficients $R[{X^{(f)}},{Y}]$ between the dynamic propensity and each structural feature $\textbf{X}^{(f)}$ of the SLO and JBB descriptors, respectively.
@@
It is instead more difficult to extract straightforward physical information from the analysis of the JBB descriptor, owing to its complex dependence on the chemical composition. It is nonetheless interesting to note that the perimeter $\overline{\pi}$ of the Voronoi cells, which is included in the JBB descriptor but not in the SLO one, has an overall negative correlation with the dynamic propensity. This again points to a coupling between steric effects, i.e., larger local free volume, and mobility.
@@
In Figures 23 and 23, we show the correlation matrices $\mathrm{C}$ of the SLO and JBB descriptors, respectively. As for the BP descriptor, the correlation matrices display a pronounced block structure: there are strong correlations within subblocks of features, this time due to similar coarse-graining lengths, as well as non-trivial correlations between different features. In particular, the SLO descriptor reveals a negative correlation between local number density $\overline{\rho}$ and local packing fraction $\overline{\varphi}$. This counter-intuitive result is likley a result of compositional fluctuations in this mixture, see Sec. III.6.
Appendix C Ordinary least squares regression and Ridge regression